OpenStudy (anonymous):

The population of a region is growing exponentially. There were 40 million people in 1990 (t = 0) and 70 million in 2000. Find an expression for the population at any time t, in years. Use the general exponential function and remember to use exact values. P(t) = (in millions) What population would you predict for the year 2010? ______ million people What is the doubling time? (Round your answer to one decimal place.) _________years

OpenStudy (anonymous):

you can use \[40\times \left(\frac{7}{4}\right)^{\frac{t}{10}}\]if you like although i doubt that is what the teacher had in mind, it is the easiest

OpenStudy (anonymous):

since 2010 is 20 years after 1990 the answer to the second question would be \[40\times \left(\frac{7}{4}\right)^{\frac{20}{10}}=40\times \left(\frac{7}{4}\right)^{2}\] \[=122.5\]

OpenStudy (anonymous):

whats the doubling time?